# CroMe: Multimodal Fake News Detection using Cross-Modal Tri-Transformer and Metric Learning

### v1

@@ 5 Conclusion
This study introduces Cross-Modal Tri-Transformer and Metric Learning (CroMe), a Multimodal Fake News Detection model. CroMe uses Bootstrapping Language-Image Pre-training with Frozen Image Encoders and Large Language Models (BLIP2) to capture intra-modality and inter-modality relationships. The model includes four main modules: encoders (BERT, BLIP2-text, Masked Autoencoders (MAE), and BLIP2-image), a metric learning module (proxy anchor method), a feature fusion module (Cross-Modal Tri-Transformer Fusion, CMTTF), and a fake news detection module. CroMe uses BLIP2 for semantic information extraction, CMTTF for feature fusion, and the proxy anchor method for metric learning. It improves accuracy by 1.3% on the Weibo dataset and 3.5% on the Weibo-21 dataset compared to previous models. CroMe performed slightly below the state-of-the-art model by 0.9% in the case of Politifact, due to the smaller dataset size.

### v2

@@ V Conclusion
This study introduces Cross-Modal Tri-Transformer and Metric Learning (CroMe), a Multimodal Fake News Detection model. CroMe uses Bootstrapping Language-Image Pre-training with Frozen Image Encoders and Large Language Models (BLIP2-OPT) to capture intra-modality and inter-modality relationships. The model includes four main modules: encoders (BERT, BLIP2-OPT-text, Masked Autoencoders (MAE), and BLIP2-OPT-image), a metric learning module (proxy anchor method), a feature fusion module (Cross-Modal Tri-Transformer Fusion, CMTTF), and a fake news detection module. CroMe uses BLIP2-OPT for semantic information extraction, CMTTF for feature fusion, and the proxy anchor method for metric learning. It improves accuracy by 1.3% on the Weibo dataset and 3.5% on the Weibo-21 dataset compared to previous models. CroMe performed slightly below the state-of-the-art model by 0.9% in the case of Politifact, due to the smaller dataset size.
[Uncaptioned image] EunJee Choi received the B.S. degree from the Department of Business Administration, Dankook University, in 2022, and the M.S. degree from the Department of Electrical and Computer Engineering, Korea University, Seoul, Republic of Korea, in 2025. Her research interests include multimodal learning, fake news detection, large language models, and visual language models.
[Uncaptioned image] Junhyun Ahn received the B.S. degree in Electronic Engineering from Korea Aerospace University, Goyang, Republic of Korea, in 2024 and is currently pursuing the M.S. degree at the Department of Electrical and Computer Engineering, Korea University, Seoul, Republic of Korea. His research interests include artificial neural networks and deep learning.
[Uncaptioned image] XinYu Piao received the B.S. degree from the School of Electrical Engineering, Korea University, in 2020, and the Ph.D. degree from the Department of Electrical and Computer Engineering, Korea University, Seoul, Republic of Korea, in 2025. He is currently a postdoctoral researcher at the Department of Electrical and Computer Engineering, Korea University, since April 2025. His research interests include systems for artificial intelligence (AI), hardware resource management, deep learning applications, and quantum computing.
[Uncaptioned image] Jong-Kook Kim (Senior Member, IEEE) received the B.S. degree in electronic engineering from Korea University, Seoul, South Korea, in 1998, and the M.S. and Ph.D. degrees from the School of Electrical and Computer Engineering, Purdue University, in 2000 and 2004, respectively. He is currently a Professor with the School of Electrical Engineering, Korea University, where he joined in 2007. He was with the Samsung SDS IT Research and Development Center from 2005 to 2007. His research interests include heterogeneous distributed computing, energy-aware computing, resource management, evolutionary heuristics, efficient systems and computing, artificial neural networks, deep learning, and systems for AI. He is a Senior Member of the IEEE and the ACM.
[…]